Grade 11Algebra

Arithmetic progrssion.
Find sum of numbers between 100 to 200 divisible by 7

adhiraj
9 Years agoGrade 11

1 Answer

Arun
9 Years ago
Indirectly we have to find the sum of 105,112,119,126........196  (because 105 will be the 1st & 196 will be the last term which are divisble by 7)
now lets find total no. of terms in this A.P.
196=105+(n-1)*7
n-1 = 91/7
n=14
now total sum S= 7[2*105 + (14-1)*7]
S= 7[210+91]
S=7*301
S=2107